OpenStudy (anonymous):

The average of two numbers is 92. The average of four different numbers is 86. What is the weighted mean of all the numbers together? 87 88 89 90

OpenStudy (anonymous):

the sum of two numbers is 92*2=184 the sum of four numbers is 86*4=344 weighed mean= (184+344)/6= 88